OpenStudy (anonymous):

what sign ( + - x / ) do i need to use for 19 3/4 [ ? ] 9 1/4 =13 1/6

OpenStudy (anonymous):

neither maths i thought at first it was - but 19 3/4 - 9 1/4 = 10 1/2

OpenStudy (anonymous):

yh thats what i thought, i spent a hour and still dont have a answer

OpenStudy (anonymous):

- leads to 10 1/2 + leads to 29 / leads to 2.135135135 x leads to 182.6875 none of the chhoices are right

OpenStudy (anonymous):

you are right - the question is wrong - maybe a typo somewhere along the way

OpenStudy (anonymous):

hmmm well, maybe they used a different fraction

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

well its a word equation

OpenStudy (anonymous):

im gonna type in the equation

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Jack spends 1 1/2 as long on his homework as jill, Last week jill spent 9 1/4 hours doing homework show answer (forget that bit the answer will be no use) the week before jack spent 19 3/5 hours doing homework. how long did jill spend doing the homework

OpenStudy (anonymous):

i answered it and it popped the answer 13 1/6 cause i done a mistake of the sign :(

OpenStudy (phi):

Jack spends 3/2 as long as jill. as an equation Jack= (3/2) Jill with Jill= 19 3/5, we have 19 3/5 = (3/2) Jill so Jill = 19 3/5 divided by 3/2 \[19\frac{3}{5}\div\frac{3}{2}= 19\frac{3}{5}\times\frac{2}{3}\] if you work it out you get 13 1/15